• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name kalyan

    Log entry time 00:09:44 on August 7,2006

    Entry number 177810

    keyword=swing shift/run summary

    Shift Summary
    --------------
    16:15 shift started with beam permit
    16:30 took the beam with empty target for checking raster size
    16:59 production started with LD2
    18.00 changed to LH2 target->production
    19:55 MCC called and took away the beam for PZT test
    20:08 beam restored to 20muA
    20:11 production started with LH2 again
    22:05 5% radiator is in and running with LH2
    22:31 Right arm run started
    23:40 4% radiator in. current changed to 12muA

    Run Summary
    ------------
    2220 empty target run for raster size check
    2221 LD2 1.95M
    2222 LD2 1.9M
    2223 LH2 junk--> NMR dipole field drift
    2224 LH2 0.58M
    2225 LH2 0.25M
    2226 LH2 0.69M
    2227 LH2 15K
    2228 LH2 50k
    2229 LH2 127K
    2230 LH2 1.3M
    2231 junk
    2232 LH2 (5% rad in) 1.1M
    2233 LH2 (5% rad in) 1.5M
    2234 LH2 (4% rad in) still running